GDP can be computed as OT-G C+I+G+NX O C+I+G+T+X O X-IM
Answer #1

Answer
option 3

GDP is a sum of final goods and services produced in an economy in a given year
The given options are GDP calculation by expenditure method as all the sum expenditure on final goods and services as well as equal to the production and it is denoted by the following identity

GDP=C+I+G+NX
where
C=consumption
I==investment spending
G=govenment spending
NX=net export =X-IM=export -import